OpenStudy (anonymous):

why is (x+deltax)^2 =x^2+2xdeltax+ deltax^2? Trying to find derivative of f(x)+ x^2 +(0.01)x +5

OpenStudy (solomonzelman):

\(\large\color{slate}{ (a+b)^2=a^2+2ab+b^2 }\)

OpenStudy (solomonzelman):

In your case: \(\large\color{slate}{ a=x }\) \(\large\color{slate}{ b=\Delta x }\)

OpenStudy (solomonzelman):

I see no purpose in this expression as far as actual problem goes, but...

OpenStudy (anonymous):

That makes sense!! Thanks! Trying to relearn calc 1 in prep for calc 2

OpenStudy (solomonzelman):

oh I didn't read the second sentence

OpenStudy (xapproachesinfinity):

i don't know why people get confused when they see that delta hehe

OpenStudy (solomonzelman):

you are trying to differentiate? I think there is likely more to this problem then...

OpenStudy (solomonzelman):

I mean more info

OpenStudy (solomonzelman):

and you mean `f(x)=` , or `f(x)+ ` ?

OpenStudy (anonymous):

f(x)=

OpenStudy (solomonzelman):

oh, I thought it is one of those limits, like Reinnmun sum formulas that you have incorrectly written... \(\large\color{slate}{ f(x)= x^2 +(0.01)x +5 }\) then you can find \(\large\color{slate}{ f'(x) }\) using simple rules.

OpenStudy (solomonzelman):

but if you are differentiating the function, \(\large\color{slate}{ f(x)= x^2 +(0.01)x +5 }\) then what does \(\large\color{slate}{ (x+\Delta x)^2 }\) have to do with anything ?

OpenStudy (anonymous):

I was just trying to simplify to find the derivative

OpenStudy (solomonzelman):

anyways, you just need to find the derivative of, \(\large\color{slate}{ f'(x)= x^2 +(0.01)x +5}\)

OpenStudy (solomonzelman):

I mean the derivative of \(\large\color{slate}{ f(x)= x^2 +(0.01)x +5}\) , excuse me

OpenStudy (solomonzelman):

apply the power rule to \(\large\color{slate}{ x^2 }\) what do you get

OpenStudy (solomonzelman):

?

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE d}{\LARGE dx}\left(\begin{matrix} x^n\\ \end{matrix}\right) =n\cdot x^{n-1} }\)

OpenStudy (solomonzelman):

(d/dx is a notation for derivative) so applying this rule, (known as the "power rule") can you finish the following? \(\large\color{black}{ \frac{\LARGE d}{\LARGE dx}\left(\begin{matrix} x^2\\ \end{matrix}\right) = }\)

OpenStudy (anonymous):

Im lost now

OpenStudy (solomonzelman):

do you know the power rule for the derivative?

OpenStudy (anonymous):

I dont remember

OpenStudy (solomonzelman):